Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability: Another Dice Question: Solving Question

Expert replies
by KevinLuk » Mon Aug 24, 2015 6:40 pm
Assume that each die has 6 sides with faces numbered 1 to 6.

What is the probability that the sum of two dice will yield a 10 or lower?


I'm having a problem with the way it is explained. Thanks.
[spoiler]

The way that I am approaching these questions is that the question is asking for probabilities. The possible outcomes (to me) in which the sums of the dice are over 10 are "5 - 6", "6 - 6", "6 - 6", and "6 - 5". Can somebody please explain to me why the second pair of 6 - 6 is not included in the OA Explanation while "5 - 6" and "6 - 5" are both considered different acceptable probabilities?

Here is the book's explanation:

Solve this problem by calculating the probability that the sum will be higher than 10, and
subtract the result from 1. There are 3 combinations of 2 dice that yield a sum higher than 10: 5 + 6,
6 + 5, and 6 + 6. Therefore, the probability that the sum will be higher than 10 is 3/36, or 1/12. The
probability that the sum will be 10 or lower is 1 - 1/12 = 11/12. [/spoiler]
Join the discussion
Source: — Problem Solving |

by DavidG@VeritasPrep » Mon Aug 24, 2015 7:06 pm
Can somebody please explain to me why the second pair of 6 - 6 is not included in the OA Explanation while "5 - 6" and "6 - 5" are both considered different acceptable probabilities?
The key is that we're talking about different dice. To conceptualize this, Imagine that the two dice are different colors, one blue and one red. The scenarios that will give us a value over 10 are:

A) Red-5/Blue-6
B) Blue-5/Red-6
C) Red-6/Blue-6

A and B are unique scenarios because we're altering which die is turning up '5' and which is turning up '6.' But clearly, Red-6/Blue-6 would be no different from Blue-6/Red-6.
Veritas Prep | GMAT Instructor

Veritas Prep Reviews
Save $100 off any live Veritas Prep GMAT Course
Join the discussion

by Max@Math Revolution » Wed Aug 26, 2015 6:47 am
Forget conventional ways of solving math questions. In PS, IVY approach is the easiest and quickest way to find the answer.


Assume that each die has 6 sides with faces numbered 1 to 6.

What is the probability that the sum of two dice will yield a 10 or lower?


I'm having a problem with the way it is explained. Thanks.

==> the probability that the sum will be smaller than, or will equal 10 = 1 - probability that the sum will be greater than 10

probability = number of cases of a specific event / total number of cases

since two dices have six numbers each, the total number of cases is 6*6=36
the case where the sum is greater than 10 is (5,6),(6,5),(6,6), 3 cases.
therefore the probability that the sum will be smaller than, or will equal 10 = 1-(3C1/36C1)=1-(3/36)=1-(1/12)=11/12



www.mathrevolution.com
l The one-and-only World’s First Variable Approach for DS and IVY Approach for PS that allow anyone to easily solve GMAT math questions.

l The easy-to-use solutions. Math skills are totally irrelevant. Forget conventional ways of solving math questions.

l The most effective time management for GMAT math to date allowing you to solve 37 questions with 10 minutes to spare

l Hitting a score of 45 is very easy and points and 49-51 is also doable.

l Unlimited Access to over 120 free video lessons at https://www.mathrevolution.com/gmat/lesson

Our advertising video at https://www.youtube.com/watch?v=R_Fki3_2vO8
Join the discussion

by [email protected] » Wed Aug 26, 2015 8:57 am
Hi KevinLuk,

You might find it beneficial to 'write everything out' - while it would take a little bit of work, you would be able to physically see everything, which would make dealing with the concept easier.

Since we are rolling two 6-sided dice, there are (6)(6) = 36 possible outcomes. Those outcomes are....

11,12,13,14,15,16
21,22,23,24,25,26
31,32,33,34,35,36
41,42,43,44,45,46
51,52,53,54,55,56
61,62,63,65,65,66

Of these 36 outcomes you can either calculate what you WANT or what you DON'T WANT (and subtract that from the number 1). The second method is easier here...

There are 3 options that DO NOT fit what you're looking for.... 3/36 = 1/12

[spoiler]1 - 1/12 = = 11/12 [/spoiler]= the options that you ARE looking for.

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion